# IDEAL: Influence-Driven Selective Annotations Empower In-Context Learners in Large Language Models

### v2

@@ D.1.7 MultiWoz
    -- Using valid SQLite, answer the following multi-turn conversational
@@ D.1.8 GeoQ
    -- Using valid SQLite, answer the following questions for the tables
@@
    -- what is the longest river in the state with the highest point
@@ D.2 Implementation details
General experimental conditions. We primarily use PyTorch (Paszke et al., 2019) to implement our algorithm and baselines. For GPT-3.5-Turbo, we perform the experiments by calling the OpenAI API using a single Intel Xeon CPU. The GPT-J 6B and GPT-Neo 2.7B models are from the Huggingface transformer library (Wolf et al., 2019). We run all our experiments of GPT-J 6B and GPT-Neo 2.7B on a single NVIDIA Tesla V100 (32GB) GPU.
Details of getting unlabeled data. Since obtaining unlabeled examples in realistic scenarios is also a high-variance process, we follow the same setting as (Su et al., 2023) to simulate the realistic setting. We perform selective annotations from 3k instances that are randomly sub-sampled from training data for each task. For each experiment, we repeat the sub-sampling process three times and average the results over all trials to ensure comprehensive evaluations.
@@
Appendix E Time Complexity Analysis
In this section, we perform a time-complexity analysis of our method. Given the constructed graph with $a$ nodes and $b$ edges, and an annotation budget of $m$, the whole algorithm involves the following two parts that incur the following time costs: (1) Information diffusion process. The time complexity of quantifying the influence of a specific subset is $\mathcal{O}(a+b)$. This is because it involves running an independent cascade diffusion process (BFS-like traversals) through the graph. (2) Greedy search. The greedy algorithm iteratively selects the example that provides the maximum marginal gain in influence. When the annotation budget is $m$, the time complexity is $\mathcal{O}(m*(a+b))$. This is because, in the worst case, the algorithm needs to evaluate the influence of each node in the network. Besides, for each node, the algorithm needs to perform a simulation with time complexity of $\mathcal{O}(a+b)$.
Appendix F Limitations
@@
The limitation of Auto-IDEAL. Auto-IDEAL outperforms IDEAL in terms of performance, but it has two main drawbacks that hinder its usability in practice. First, Auto-IDEAL suffers from a model inference cost, especially in the era of large language models as a service. Specifically, when performing automatic annotation, Auto-IDEAL has to make predictions for all unlabeled examples to achieve automatic annotation. This makes it less practical than the native IDEAL. Second, Auto-IDEAL may fail when the initial examples to be labeled are not very relevant to the initial examples labeled, even though they have a similar embedding. Auto-IDEAL performs automatic annotation by following the information diffusion process from the initial annotated subset to the examples to be labeled with similar embedding. When the examples to be labeled are not relevant to the initial examples labeled, it may lead to incorrect automatic annotations and then poor performance. Future research may focus on maintaining superior performance while reducing the automatic annotation cost of IDEAL.
Potential benchmark leakage in GPT-3.5-Turbo. There may potentially exist benchmark leakage problems in the evaluation process (Zhou et al., 2023). Specifically, due to the fact that GPT-3.5-Turbo is trained using huge text datasets as of 2021, the data related to evaluation sets may potentially be used for model training. This could lead to potential risks in the evaluation process. However, as our work does not involve the training data selection, the impact should be negligible.

### v3

@@ D.1.7 MultiWoz
-- Using valid SQLite, answer the following multi-turn conversational
@@ D.1.8 GeoQ
-- Using valid SQLite, answer the following questions for the tables
@@
-- what is the longest river in the state with the highest point
@@ D.2 Implementation details
General experimental conditions. We primarily use PyTorch [79] to implement our algorithm and baselines. For GPT-3.5-Turbo, we perform the experiments by calling the OpenAI API using a single Intel Xeon CPU. The GPT-J 6B and GPT-Neo 2.7B models are from the Huggingface transformer library [16]. We run all our experiments of GPT-J 6B and GPT-Neo 2.7B on a single NVIDIA Tesla V100 (32GB) GPU.
Details of getting unlabeled data. Since obtaining unlabeled examples in realistic scenarios is also a high-variance process, we follow the same setting as [7] to simulate the realistic setting. We perform selective annotations from 3k instances that are randomly sub-sampled from training data for each task. For each experiment, we repeat the sub-sampling process three times and average the results over all trials to ensure comprehensive evaluations.
@@
Appendix E Limitations
@@
Time cost of Auto-IDEAL. Although Auto-IDEAL achieves even better performance than IDEAL, it has the same drawback as Vote-$k$. That is to say, when making automatic annotations, it incurs the cost of making predictions for all unlabeled data. Future work may study how to maintain superior performance while reducing the automatic annotation cost of IDEAL at the same time. Compared with [7], we do not evaluate NQ [80] due to budget constraints.